OpenStudy (anonymous):

the number of men in a tour group is 15 times the number of woman if there are 48 people in the group how many are men ?

OpenStudy (texaschic101):

M + W = 48 M = 15W Now sub 15W + W = 48 16W = 48 W = 48/16 W = 3 M = 15W M = 15(3) M = 45 There are 45 men (M) and 3 women (W)